OpenStudy (anonymous):

Solve -2(2x + 5) - 3 = -3(x - 1).

Mehek (mehek14):

distribute \(-4x-10-3=-3x+3\) \(-4x-13=-3x+3 \\ -4x = -3x + 16 \\ -x = 16 \\ x = -16\)

Mehek (mehek14):

if you plug -16 into the equation you get \(51=51\)

OpenStudy (anonymous):

o okay i understand thanks

Mehek (mehek14):

yw